Get Help

Online: Request Help
Phone: (707) 826-HELP (4357)
Email: help@humboldt [dot] edu
In Person: Library 120 • Hours
Reset HSU Password
System Status

Services For

System Status Archive

System Issue start Who is affected What happened
CMS/Peoplesoft Sep 8 2011 - 18:00 CFS (Finance) or CFS Data Warehouse users and programmers CMS Central will apply a campus interface customization. Refer to CO's CFSPRD 1048 2011 Bulletin dated September 6, 2011.
Network Folders Sep 1 2011 - 12:00 All Network Folders Users The Winbind process stopped working, and had to be restarted. It has since been set to restart at night Mondays, and Thursdays.
Wireless Network Sep 1 2011 - 05:00 Any user accessing the HSUWireless, ResnetWireless or Resnet Wired Network In order to resolve an important security issue, the TCP port used in the Safe Connect messaging portal at https://nac.humboldt.edu:8443/ will be changed from 8443 to 9443 or https://nac.humboldt.edu:9443/ . Be sure to grant the Safe Connect process access to port 9443 through your firewall if prompted. If you are repeatedly prompted to install the Policy Key, even though it is running - check your firewall settings before contacting the Help Desk or ResNet Desk if you live on campus.
Campus Phone System Aug 26 2011 - 09:15 All campus users trying to reach voice mail, some callers trying to reach some campus extensions Problem has resolved. Cause is still being investigated
Campus Data Network Aug 25 2011 - 03:00 All network users in all Houses, Marine Lab at Trinidad, Natural History Museum, University Center, and Resnet, and all Wireless users Due to De-gradated service, TNS will be restarting the Siemens Hall Core Switch. If your network service does not come up after that time frame, reboot your system.
CMS/Peoplesoft Aug 25 2011 - 18:00 Anyone who uses PeopleSoft Finance (CFS) CMS Central will install fixes that do not affect campus users, except that the system will not be available. Reference CFS Bulleting 1045 2011
Wireless Network Aug 22 2011 - 11:35 Campus wireless users Allocated IP space is exhausted. Resolution: Allocate additional IP space
CMS/Peoplesoft Sep 4 2011 - 00:01 Anyone who uses PeopleSoft or runs CRON jobs from 9/3 at 5 p.m.-9/4 6:30 a.m. Chancellor's Office ITS Support Center monthly maintenance. Reference bulletin ITS 2011 0170
Other Aug 19 2011 - 17:30 All Users of SHC-Term, SHC-Pyramed More RAM will be added, systems need to be restarted.
OBI Aug 16 2011 - 09:56 Anyone trying to use OBI reporting system Issues with an incomplete change over to the new fiscal year and the unanticipated reuse of an inactive department code caused parts of the ETL process to fail.
Other Aug 16 2011 - 16:00 All Webcheckout Users Upgrade from this a.m. had to be postponed by Onshore.
Other Aug 16 2011 - 13:00 All Webcheckout Users The software will be upgraded to a newer version.
Pharos Pay-for-Print Aug 12 2011 - 15:00 Anyone printing in computer labs. We are updating the script to allow for grant billing and wireless printing. The script has been tested in development and is being put into production during this time. The change should be transparent to people printing in the labs. The script has been changed and is working as expected.
CMS/Peoplesoft Aug 21 2011 - 00:01 Finance (CFS) and anyone else who uses the CSU Portal CO Infrastructure Services will be performing scheduled maintenance (reference CO ITS Support Center bulletin 2011 0160)
Account Center Aug 10 2011 - 10:00 Account Center users Account Center went down for a few minutes, but is now back up.
OBI Aug 9 2011 - 17:00 OBI Report system users We'll be upgrading the OBI 11.1.1.3 version to OBI 11.1.1.5
CMS/Peoplesoft Aug 11 2011 - 18:00 Finance (CFS and CFS Data Warehouse) Per Chancellor's Office CFSPRD 1043 2011 bulletin: CFS Production (FCFSPRD) and CFS Data Warehouse application and database services will be unavailable, including process schedulers and integration broker. End User functionality within the CFS application will not change with this update.
Moodle Aug 8 2011 - 15:45 Faculty editing course sections An error in an update to the course sections caused a problem with updates made today. We are restoring from this morning's backup to resolve the problem.
Other Aug 14 2011 - 05:00 All Users Data Center Systems will be patched, and rebooted. The date was moved forward by a week.
OBI Aug 7 2011 - 09:00 Anyone using OBI.
Campus Data Network Aug 5 2011 - 14:45 All Users trying to access Off-Campus Sites and email The Campus has lost access to the internet. The problem is outside of HSU. We are looking into the problem. There is no estimate at this time. Please check back for updates. This was fixed late friday evening.
Off-Campus Internet Connection Aug 5 2011 - 14:40 All persons using the network to access off campus Cenic reports that data circuits into Humboldt County are down. They are investigating. Check back for updates. This was fixed late friday evening.
OBI Aug 2 2011 - 13:18 OBI Report system users The service was restarted because of poor performance.
Pharos Pay-for-Print Aug 2 2011 - 17:30 Users printing in Labs A new license file will be installed on the Pharos Pay4Print servers.
Other Aug 2 2011 - 17:00 SHC users of Titanium and MedPro servers Network connections will be migrated to different jacks in the Data Center.
Campus Data Network Jul 30 2011 - 04:00 All users inside SBS building 100 excepting UPD The Telecom Room in SBS has been notoriously hot, and the excess heat over the past six years has taken a toll on some of the switching equipment, which is exhibiting symptoms of premature failure. Network performance in the building will be noticeably improved using the new switching modules. HVAC installation is also in the planning stages.
Campus Data Network Jul 30 2011 - 04:00 All users inside SBS Building 100 (except UPD) The Telecom Room in SBS has been notoriously hot, and the excess heat over the past six years has taken a toll on some of the switching equipment, which is exhibiting symptoms of premature failure. Network performance in the building will be noticeably improved using the new switching modules. HVAC installation is also in the planning stages.
Directory Services Jul 23 2011 - 11:00 People attempting to login to Google and CFS Software configuration patch to address security issues.
Other Jul 20 2011 - 20:00 All users Roles are be changed and some service pack changes will be done
CMS/Peoplesoft Jul 20 2011 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions HCM 8.9 Maintenance Pack 13.01 and Student Services Summary modification update will be installed by CMS Central this evening.
feedback